OpenStudy (anonymous):

-1 in a square root sign with the negative outside

myininaya (myininaya):

\[-\sqrt{-1}=-(i)=-i\]

OpenStudy (anonymous):

no the one doesn't hav a negative inside the square root

myininaya (myininaya):

\[-\sqrt{1}=-(1)=-1\]

OpenStudy (anonymous):

really thats it?

myininaya (myininaya):

yep lol

OpenStudy (anonymous):

wow im dumb

OpenStudy (anonymous):

also \[\sqrt{3} - 7\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{3}(1-7)=?\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

1-7=?

OpenStudy (anonymous):

answer it

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

so \[-6\sqrt{3}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

also \[2\sqrt{5}-3\sqrt{45}\]

OpenStudy (phi):

rewrite the sqrt(45) as sqrt(9*5) and go from there

OpenStudy (anonymous):

\[2\sqrt{5}-3\sqrt{5*9}=2\sqrt{5}-9\sqrt{5}=-7\sqrt{5}\]

OpenStudy (anonymous):

wouldnt the 9 be 3 cause three is the sqrt of 9

OpenStudy (anonymous):

makin it -1sqrt5

OpenStudy (anonymous):

\[-1\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{9}=3\] \[3*3=9\]

OpenStudy (anonymous):

where did the second three come from?

myininaya (myininaya):

sorry chels i was doing another problem

OpenStudy (anonymous):

its ok

myininaya (myininaya):

hes just showing you why \[\sqrt{9}=3\]

OpenStudy (anonymous):

ok so the answer is\[-7\sqrt{5}\]?

myininaya (myininaya):

yes

OpenStudy (anonymous):

ok thanks everyone on this